OpenStudy (anonymous):

what is 20 percent of 300?

OpenStudy (lalaly):

300*0.2 = 60

OpenStudy (anonymous):

20 percent can also be seen as one-fifth, dividing 300 by 5 gives you the same answer.

OpenStudy (anonymous):

hmmm..........who is right here?

OpenStudy (anonymous):

We are both right, it's just two different methods. 300/5=60, same as lalaly!

OpenStudy (anonymous):

oh i thought you said multiply 300 by 5 haha(: sorry

OpenStudy (anonymous):

so you can go different way. i thik the easy way 300 = 100 % 300 /100 = 3 so 3 = 1% 3 * 20% = 60 so 60 = 20% ;-) or you can get this formular 100% = 300 20% = x calc 300*20% / 100% = 60